OpenStudy (anonymous):

If Log 17 = K determine an expression for the following: Log 170

OpenStudy (anonymous):

this means: 10^k=170

OpenStudy (anonymous):

K with a 0 at the end lol idk sry

OpenStudy (anonymous):

an example in class was done like this: Log1700 Log 17*100 Log 17 + Log 100 K + Log_{10} 10^2 =K+2 but when I did the same thing for the above question the answer was wrong according to the book....

OpenStudy (anonymous):

ooh ok i see what is going on: so if log 17 =k then log 170 = log 17*10 = log 17 + log 10 = k + log 10 = K + 1

OpenStudy (anonymous):

ya thats what I got...Maybe my textbook is wrong lol

OpenStudy (anonymous):

are you sure you wrote it down correctly?

OpenStudy (anonymous):

theres a screenshot